OpenStudy (anonymous):

13. A tree casts a shadow of 30ft. The angle of elevation from the tip of the shadow to the top of the tree is 70 degrees. To the nearest foot, what is the height of the tree?

OpenStudy (anonymous):

827

OpenStudy (johnweldon1993):

 you have an angle...a side adjacent to that angle...and you want the side opposite that angle. so which trig function uses Opposite and Adjacent?

OpenStudy (anonymous):

@johnweldon1993 I think cos but I'm probably wrong haha

OpenStudy (johnweldon1993):

to remember the trig functions...remember SOH CAH TOA Sin = Opposite / Hypotenuse Cos = Adjacent / Hypotenuse Tan = Opposite / Adjacent so which is it...?

OpenStudy (anonymous):

Tan

OpenStudy (johnweldon1993):

Great....it is the tangent function So \[\tan = \frac{ Opposite }{ Adjacent }\] You would plug in the values you have \[\tan 70 = \frac{ Opposite }{ 30 }\] and you want to solve for opposite...so what do you do...?

OpenStudy (anonymous):

Uhm.. I'm not sure :/

OpenStudy (johnweldon1993):

No problem.....you would multiply both sides by 30 right? when you multiply \[\frac{ Opposite }{ 30 } * 30\] the 30's cancel out so now you would have \[Opposite = \tan 70 * 30\] So...what is the tan of 70 degrees?

OpenStudy (anonymous):

I got 82.42 and a bunch of other numbers after that

OpenStudy (johnweldon1993):

ahhh skipped ahead a step :) lol yes....now the question asks you "to the nearest foot" so what is 82.42 rounded to the nearest whole number?

OpenStudy (anonymous):

Well I don't think I can round up so it would just be 82, right?

OpenStudy (anonymous):

Or wait, maybe it could be 83

OpenStudy (johnweldon1993):

That is correct....*remember 82.49999999999999999999999999 is rounded to 82.... as soon as it hits 82.5...you would round up so here you only have 82.42...so it would be...?

OpenStudy (anonymous):

It would just be 82 :) I knew that you can only round up if the it's a 5 or more

OpenStudy (johnweldon1993):

There you go :)

OpenStudy (anonymous):

Thank you! :)

OpenStudy (johnweldon1993):

anytime!